OpenStudy (anonymous):

-3(5 - 15n) = 15(3n - 2)

OpenStudy (anonymous):

-3(5-15n)=15(3n-2) -15+45n=45n-30 45n will be cancelled -15=-30 1=2

OpenStudy (campbell_st):

This equation doesn't have a solution

OpenStudy (anonymous):

thanks! how about this one: 1/4x - 5 = x

OpenStudy (anonymous):

So the equatin has no solution

OpenStudy (campbell_st):

3/4 x = -5 3x = -20 x = -20/3

OpenStudy (anonymous):

2/3y + 6 = y

OpenStudy (anonymous):

45 - 4r = 45 -4r

OpenStudy (anonymous):

2/3y+6=y 2+3y(6)/3y=y 2+18y/3y=y 2+18y=y3y^2 3y^2-18y=2

OpenStudy (anonymous):

whats the answer?

OpenStudy (campbell_st):

y = 18